OpenStudy (anonymous):

for the function f, costruct and simplify the difference quotient f(x+h)-(x) all over h. f(x)=x/7-x

OpenStudy (anonymous):

this requires a raft of algebra

OpenStudy (anonymous):

\[\frac{\frac{x+h}{7-(x+h)}-\frac{x}{7-x}}{h}\] is a start